# Streaming algorithms for products of probabilities

### v1

@@ 4 The threshold problem
Take the set of words $S_{n,b}$, the numbers $s_{i}$ and the threshold $t$ from the proof of Theorem 4.3. Let $f:\{1,2,\ldots,B^{n}\}\to S_{n,b}$ be the unique bijection such that for all $1\leq i,j\leq B^{n}$ we have: $i<j$ if and only if the product of the numbers in the word $f(i)$ is strictly smaller than the product of the numbers in the word $f(j)$.
On input $i$, Alice computes the word $f(i)=q_{1}q_{2}\cdots q_{n}$, reads $t\,q_{1}\cdots q_{n}$ into $\mathcal{R}$ and sends the resulting memory state $\alpha$ of $\mathcal{R}$ to Bob. Bob then computes from his input $j$ the word $f(j)=r_{1}r_{2}\cdots r_{n}$ and reads, starting from memory state $\alpha$, the word $(s_{1}/r_{1})(s_{2}/r_{2})\cdots(s_{n}/r_{n})$ into the algorithm. From the final output of $\mathcal{R}$, Bob can decide with probability at least $2/3$ whether $i<j$:
  • If $\prod_{i=1}^{n}q_{i}<\prod_{i=1}^{n}r_{i}$ (i.e., $i<j$) then $\mathcal{R}$ outputs $1$ with probability at least $2/3$ because
    $\prod_{i=1}^{n}q_{i}\prod_{i=1}^{n}\frac{s_{i}}{r_{i}}<\prod_{i=1}^{n}s_{i}=t.$
  • If $\prod_{i=1}^{n}q_{i}\geq\prod_{i=1}^{n}r_{i}$ (i.e., $i\geq j$) then $\mathcal{R}$ outputs 00 with probability at least $2/3$ because
    $\prod_{i=1}^{n}q_{i}\prod_{i=1}^{n}\frac{s_{i}}{r_{i}}\geq\prod_{i=1}^{n}s_{i}=t.$
It follows that the communicated memory state $\alpha$ of $\mathcal{R}$ must have bit length $\Omega(\log B^{n})=\Omega(n\cdot b)$.
Corollary 4.7.
Every randomized streaming algorithm $\mathcal{R}$ for $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{R}}(n,b)\geq\Omega(n\cdot b)$ whenever $b\geq\Omega(\log_{2}n)$.
The condition $b\geq\Omega(\log_{2}n)$ cannot be completely avoided in Corollary 4.7, since this would contradict the upper bound $\mathcal{O}(2^{b})$ in Theorem 4.1.
Also note that the lower bound for $\mathsf{APP}_{\epsilon}$ in Theorem 3.2 holds for input streams over a set of only 3 numbers (namely the set $S$ in (5)). The lower bound from Corollary 4.7 does not hold for a constant number of probabilities, since we can then store the product with $\mathcal{O}(\log n)$ bits.
@@ 5 Sliding window products
Theorem 5.1.
For $20<\epsilon<1/$ 2, there are deterministic sliding window algorithms for $\mathsf{SWAPP}_{m,\epsilon}$ with space complexity $m\cdot(\log m+\log b-\log\epsilon)$, respectively $2mb$.
Proof 5.2.
The upper bound $2mb$ is clear, since the whole window content can be stored with $2mb$ bits. For the upper bound $m\cdot(\log m+\log b-\log\epsilon)$ we use the algorithm for $\mathsf{APP}_{\epsilon}$ from [1]. Recall the sketch after Theorem 3.1. Our sliding window algorithm stores for a window content $q_{1}q_{2}\cdots q_{m}$ the sequence of bucket indices $a_{1},a_{2},\ldots,a_{m}$, where $q_{i}\in B_{a_{i},\epsilon^{\prime}}$ for $\epsilon^{\prime}=\epsilon/m$. By the analysis in [1], $(1-\epsilon^{\prime})^{a}$ (with $a=\sum_{i=1}^{m}a_{i})$ is a $(1-\epsilon)$-approximation of $\prod_{i=1}^{m}q_{i}$. Moreover, every $a_{i}$ is bounded by $mb/\epsilon$, which yields the space bound from the theorem.
Theorem 5.3.
Fix an $\epsilon\in(0,1/2)$. Every randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ requires space $\Omega(m\cdot(\log b-\log\epsilon-\Theta(1)))$.
Proof 5.4.
@@
Define the following additional integers:
@@
Consider the set $A=\{2^{-i\cdot\alpha}:0\leq i\leq c-1\}\subseteq\mathbb{P}$ of size $c$. Also note that $\|q\|\leq c\cdot\alpha=b$ for all $q\in A$. We prove the theorem by a reduction to the communication problem $\mathsf{IndexGT}_{m,n}$.
Assume that we have a randomized sliding window algorithm $\mathcal{R}$ for $\mathsf{SWAPP}_{\epsilon}$. We can assume that the error probability of $\mathcal{R}$ is bounded by $1/6$. We obtain a randomized one-way communication protocol for $\mathsf{IndexGT}_{c,m}$ as follows: The input of Alice is a sequence $a_{1},a_{2},\ldots,a_{m}$ with $a_{i}\in A$ and the input of Bob is an index $1\leq i\leq m$ and a number $a\in A$. He wants to find out whether $a_{i}<a$. Alice sends the memory state of the randomized sliding window algorithm after reading $a_{1}a_{2}\cdots a_{m}$ to Bob. This allows Bob to compute with high probability $(1-\epsilon)$-approximations $P^{\prime}_{1}$ and $P^{\prime}_{2}$ of the products
$P_{1}=a_{i}a_{i+1}\cdots a_{m}\quad\text{ and }\quad P_{2}=a_{i+1}\cdots a_{m}a.$
For this, Bob continues the sliding window algorithm from the memory state obtained from Alice with the input numbers $1,1,\ldots,1$ and $1,1,\ldots,1,a$, respectively, where the number of $1$’s is $i-1$. With probability at least $1-1/3$ the computed approximations $P^{\prime}_{1}$ and $P^{\prime}_{2}$ satisfy
$P^{\prime}_{1}\in((1-\epsilon)P_{1},P_{1}/(1-\epsilon))\quad\text{ and }\quad P^{\prime}_{1}\in((1-\epsilon)P_{2},P_{2}/(1-\epsilon)).$ (10)
Assume for the following that (10) holds. We claim that from the quotient $P^{\prime}_{1}/P^{\prime}_{2}$, Bob can determine whether $a_{i}<a$ holds. Note that $P_{1}/P_{2}=a_{i}/a$. Moreover, (10) implies
$\frac{P^{\prime}_{1}}{P^{\prime}_{2}}\in\bigg{(}(1-\epsilon)^{2}\cdot\frac{a_{i}}{a},\;\frac{1}{(1-\epsilon)^{2}}\cdot\frac{a_{i}}{a}\bigg{)}.$
If $a_{i}\leq a$ then $P^{\prime}_{1}/P^{\prime}_{2}<1/(1-\epsilon)^{2}$. On the other hand, if $a_{i}>a$ then by the choice of the set $B$ we must have $a_{i}/a\geq 2^{\alpha}\geq 2^{4\delta}=1/(1-\epsilon)^{4}$, which yields $P^{\prime}_{1}/P^{\prime}_{2}>1/(1-\epsilon)^{2}$. Hence, from the quotient $P^{\prime}_{1}/P^{\prime}_{2}$, Bob can indeed distinguish the cases $a_{i}>a$ and $a_{i}\leq a$.
By Theorem 2.1, every randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ must use space $\Omega(m\cdot\log c)=\Omega(m\cdot(\log b-\log\delta-\Theta(1)))$.
6 Open problem
For the space complexity of $\mathsf{SWAPP}_{m,\epsilon}$, there is a gap between the upper bounds in Theorem 5.1 and the lower bound from Theorem 5.3 that we would like to close. For $\mathsf{APP}_{\epsilon}$, a smaller gap (by a multiplicative factor of 2) arises from the upper bound in Theorem 3.1 and the lower bound in Theorem 3.2. One may also ask, whether the condition $-\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$ in Theorem 3.2 can be relaxed.

### v2

@@ 4 The threshold problem
Take the set of words $S_{n,b}$ from (10), the numbers $s_{i}$ from (12) and the threshold $t$ from (11). Let $f:\{1,2,\ldots,B^{n}\}\to S_{n,b}$ be the unique bijection such that for all $1\leq i,j\leq B^{n}$ with $f(i)=q_{1}q_{2}\cdots q_{n}$ and $f(j)=r_{1}r_{2}\cdots r_{n}$ we have: $i>j$ if and only if $\prod_{k=1}^{n}q_{k}<\prod_{k=1}^{n}r_{k}$.
On input $i$, Alice computes the word $f(i)=q_{1}q_{2}\cdots q_{n}$, reads $t\,q_{1}\cdots q_{n}$ into $\mathcal{R}$ and sends the resulting memory state $\alpha$ of $\mathcal{R}$ to Bob. Bob then computes from his input $j$ the word $f(j)=r_{1}r_{2}\cdots r_{n}$ and reads, starting from memory state $\alpha$, the word $(s_{1}/r_{1})(s_{2}/r_{2})\cdots(s_{n}/r_{n})$ into the algorithm. From the final output of $\mathcal{R}$, Bob can decide with probability at least $2/3$ whether $i>j$:
  • If $\prod_{k=1}^{n}q_{k}<\prod_{k=1}^{n}r_{k}$ (i.e., $i>j$) then $\mathcal{R}$ outputs $1$ with probability at least $2/3$ because
    $\prod_{k=1}^{n}q_{k}\prod_{k=1}^{n}\frac{s_{k}}{r_{k}}<\prod_{k=1}^{n}s_{k}=t.$
  • If $\prod_{k=1}^{n}q_{k}\geq\prod_{k=1}^{n}r_{k}$ (i.e., $i\leq j$) then $\mathcal{R}$ outputs 0 with probability at least $2/3$ because
    $\prod_{k=1}^{n}q_{k}\prod_{k=1}^{n}\frac{s_{k}}{r_{k}}\geq\prod_{k=1}^{n}s_{k}=t.$
It follows from Theorem 2.1 that the communicated memory state $\alpha$ of $\mathcal{R}$ must have bit length $\Omega(\log B^{n})=\Omega(n\cdot b)$. ∎
If we require $b\geq\Omega(\log_{2}n)$ then the term $(2+\gamma)(\log_{2}n+b)$ in Theorem 4.3 becomes $\Theta(b)$ and we obtain:
Corollary 1
Every randomized streaming algorithm $\mathcal{R}$ for the threshold problem $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{R}}(n,b)\geq\Omega(n\cdot b)$ whenever $b\geq\Omega(\log_{2}n)$.
The condition $b\geq\Omega(\log_{2}n)$ cannot be completely avoided in Corollary 1, since this would contradict the upper bound $1.443\cdot 2^{b}$ in Theorem 4.1.
Also note that the lower bound for $\mathsf{APP}_{\epsilon}$ in Theorem 3.2 holds for input streams over a set of only 3 numbers (namely the set $S$ in (6)). The lower bound from Corollary 1 does not hold for a constant number of probabilities, since we can then store the product with $\mathcal{O}(\log n)$ bits.
@@ 5 Sliding window products
Theorem 5.1
For $0<\epsilon<1/2$, there are deterministic sliding window algorithms for $\mathsf{SWAPP}_{m,\epsilon}$ with space complexity $m\cdot(\log m+\log b-\log\epsilon)$ and $2mb$, respectively.
Proof
The upper bound $2mb$ is clear, since the whole window content can be stored with $2mb$ bits. For the upper bound $m\cdot(\log m+\log b-\log\epsilon)$ we use the algorithm for $\mathsf{APP}_{\epsilon}$ from [2]. Recall the sketch after Theorem 3.1. Our sliding window algorithm stores for a window content $q_{1}q_{2}\cdots q_{m}$ the sequence of bucket indices $a_{1}a_{2}\cdots a_{m}$, where $q_{i}\in B_{a_{i},\epsilon^{\prime}}$ for $\epsilon^{\prime}=\epsilon/m$. By the analysis in [2], $(1-\epsilon^{\prime})^{a}$ (with $a=\sum_{i=1}^{m}a_{i})$ is a $(1-\epsilon)$-approximation of $\prod_{i=1}^{m}q_{i}$. Moreover, every $a_{i}$ is bounded by $mb/\epsilon$ and therefore can be stored with $\log m+\log b-\log\epsilon$ bits. ∎
Theorem 5.2
Let $0<\epsilon<1/2$. Every randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ requires space $\Omega(m\cdot(\log b-\log\epsilon-\Theta(1)))$.
Proof
@@
Define the following integers:
@@
Consider the set $A=\{2^{-i\cdot\alpha}:0\leq i\leq c-1\}\subseteq\mathbb{P}$ of size $c$. Also note that $\|q\|\leq c\cdot\alpha\leq b$ for all $q\in A$. We prove the theorem by a reduction from the communication problem $\mathsf{IGT}_{c,m}$.
Let $\mathcal{R}$ be a randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$. We can assume that the error probability of $\mathcal{R}$ is bounded by $1/6$. We obtain a randomized one-way communication protocol for $\mathsf{IGT}_{c,m}$ as follows: The input of Alice is a sequence $a_{1},a_{2},\ldots,a_{m}$ with $a_{i}\in A$ and the input of Bob is an index $1\leq i\leq m$ and a number $a\in A$. Bob wants to find out whether $a_{i}>a$. Alice sends the memory state of the randomized sliding window algorithm after reading $a_{1}a_{2}\cdots a_{m}$ to Bob. This allows Bob to compute with high probability $(1-\epsilon)$-approximations $P^{\prime}_{1}$ and $P^{\prime}_{2}$ of the products
$P_{1}=a_{i}a_{i+1}\cdots a_{m}\quad\text{ and }\quad P_{2}=a_{i+1}\cdots a_{m}a,$
respectively. For this, Bob continues the sliding window algorithm from the memory state obtained from Alice with the input numbers $1,1,\ldots,1$ and $1,1,\ldots,1,a$, respectively, where the number of $1$’s is i1i-1. With probability at least 11/31-1/3 the computed approximations $P^{\prime}_{1}$ and $P^{\prime}_{2}$ satisfy
$(1-\epsilon)P_{1}<P^{\prime}_{1}<P_{1}/(1-\epsilon)\quad\text{ and }\quad(1-\epsilon)P_{2}<P^{\prime}_{2}<P_{2}/(1-\epsilon).$ (16)
Assume for the following that (16) holds. We claim that from the quotient $P^{\prime}_{1}/P^{\prime}_{2}$, Bob can determine whether $a_{i}>a$ holds. Note that $P_{1}/P_{2}=a_{i}/a$. Moreover, (16) implies
$(1-\epsilon)^{2}\cdot\frac{a_{i}}{a}\ <\ \frac{P^{\prime}_{1}}{P^{\prime}_{2}}\ <\ \frac{1}{(1-\epsilon)^{2}}\cdot\frac{a_{i}}{a}.$
If $a_{i}\leq a$ then $P^{\prime}_{1}/P^{\prime}_{2}<1/(1-\epsilon)^{2}$. On the other hand, if $a_{i}>a$ then by the choice of the set $A$ we must have $a_{i}/a\geq 2^{\alpha}\geq 2^{4\delta}=1/(1-\epsilon)^{4}$, which yields $P^{\prime}_{1}/P^{\prime}_{2}>1/(1-\epsilon)^{2}$. Hence, from the quotient $P^{\prime}_{1}/P^{\prime}_{2}$ Bob can indeed distinguish the cases $a_{i}>a$ and $a_{i}\leq a$. By Theorem 2.1, $\mathcal{R}$ must use space $\Omega(m\cdot\log c)=\Omega(m\cdot(\log b-\log\delta-\Theta(1)))$. ∎
6 Open problems
For the space complexity of $\mathsf{SWAPP}_{m,\epsilon}$, there is a gap between the upper bounds in Theorem 5.1 and the lower bound from Theorem 5.2 that we would like to close. For $\mathsf{APP}_{\epsilon}$, a small gap (by a multiplicative factor of 2) arises from the upper bound in Theorem 3.1 and the lower bound in Theorem 3.2. One may also ask, whether the condition $\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$ in Theorem 3.2 can be relaxed.